OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: x – 5y > –7 3x + 4y –8

OpenStudy (kainui):

The second one isn't an inequality!

OpenStudy (anonymous):

its an expression

OpenStudy (anonymous):

i mean < with the line under

OpenStudy (anonymous):

its like <_

OpenStudy (slaaibak):

\[\le\]

OpenStudy (anonymous):

ya but its only 1 inequality right?

OpenStudy (kainui):

\[x-5y>-7\]\[3x + 4y \le–8\] Is this your question?

OpenStudy (anonymous):

YES THATS MY QUESTION

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I DONT REMEMBER THIS

OpenStudy (anonymous):

would it be x<-8/9 y≀2/9

OpenStudy (anonymous):

doh every 1 went away

OpenStudy (anonymous):

wait lemme check myself ok xD

OpenStudy (anonymous):

HAH OKAY:)

OpenStudy (anonymous):

THERE SUPPOSE TO BE POINTS

OpenStudy (anonymous):

a (–1, –1) b (–2, 4) c (3, –8) d (4, 5)

OpenStudy (anonymous):

lol good thing i waitedi was wrong :D

OpenStudy (anonymous):

can you help me though?

OpenStudy (anonymous):

everytime i think its a race to get the answer thats y :P

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you mean on a coordinate plane?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

those are the answers i put up top i just cant figure it out ive been on it for like 17 minutes lol

OpenStudy (anonymous):

okies much different anser than

OpenStudy (anonymous):

then you would also have to specify graphing also :P

OpenStudy (anonymous):

do you know how to put things in slop intercept form?

OpenStudy (anonymous):

i forgot

OpenStudy (anonymous):

is y=mx+b ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do so to the equations (inequalities)

OpenStudy (anonymous):

HUH? LOL

OpenStudy (anonymous):

did i say something wrong again? :P

OpenStudy (anonymous):

inequalities are almost the same as equations just that they have the <, > sign

OpenStudy (anonymous):

when you divide, multiply by a negative number the signs < (for example) switches to >

OpenStudy (anonymous):

YEAH YOU LOST ME LOL I HAVE A BRAIN FART

OpenStudy (anonymous):

for the first 1x-5y>-7 it would equal ?

OpenStudy (anonymous):

so you try :D

OpenStudy (anonymous):

i mean to put it in slope intercept form what would it look like

OpenStudy (anonymous):

ahh idk i give up lol

OpenStudy (anonymous):

okies slope intercept form as you said looks somewhat like this y=mx+b (which is correct) right?

OpenStudy (anonymous):

and you have x-5y > -7

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

umm do you know how to put simple equations into slope intercept form?

OpenStudy (anonymous):

i dont remember i prob do but i dont remember lol and i have 2 minutes left on my time to answer the QUESTION!

OpenStudy (anonymous):

y <1/5x+7

OpenStudy (anonymous):

that doesnt match any of the answers though

OpenStudy (anonymous):

y≀-3x/4-2

OpenStudy (anonymous):

(–1, –1) (–2, 4) (3, –8) (4, 5)

OpenStudy (anonymous):

if you know how to graph it should be very easy

OpenStudy (anonymous):

the time passed its too late anyways lol

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh and what was the real answer anyway haha ?

OpenStudy (anonymous):

inequalities cover a range really :P

OpenStudy (anonymous):

i dont really know about a specific point

OpenStudy (anonymous):

unless its from the point you are graphing from

OpenStudy (anonymous):

you know how to graph inequalities right?